# Mini-Omni2: Towards Open-source GPT-4o Model with Vision, Speech and Duplex

### v1

@@ 3.3 Training Strategies
  • Post training In Stage 3, the task of the model is to extend the output modality to include audio response generation. As shown in Figure 5, the model will be trained on all tasks from Stage 1 and Stage 2, with audio token outputs for all question-answering tasks. Additionally, the model will learn semantic interruption, an algorithm introduced in the next section.
3.4 Semantic Interruption
We believe that a real-time conversation model needs to be interruptible by humans in order to enable more flexible interactions. However, this interruption mechanism should not be a simple VAD (Voice Activity Detection)-based one, but rather a system that can determine whether the user intends to interrupt the model. Additionally, the model’s training in this aspect should be highly robust, capable of handling various external situations (e.g., noise, other conversations, and unrelated sounds).
Background Noise Selection: (1) We randomly utilized a variety of speech recognition samples from the Libri-tts dataset as the original human noise data samples. (2) We employed samples from the MUSAN dataset, which includes music, human voices, white noise, and urban noise.
@@ 3.4 Semantic Interruption
Combining the aforementioned data, the model will receive long sequences of data containing "Stop Omni" phrases amidst various noises. The model will generate two types of state tokens in real time: irq and n-irq, representing the intention of the user to interrupt and not to interrupt, respectively. Each time the model generates an audio token, it will predict a new task; when the irq token is output, it will interrupt the audio stream writing process. For this task, we use tokens as input to enhance the model’s real-time processing capabilities. We observed the interruption mechanism in Vita and proposed a method for integrating new tasks in batch form, allowing a unified model to control itself in real time, thereby reducing memory overhead by half compared to a dual-model approach.
@@ 4 Data and Evaluation
In this section, we introduce the data used for training Mini-Omni2 and present some initial evaluation results. We will provide a more detailed explanation of the composition and construction process of the data for each modality. In the experimental results section, we only showcase a few application cases and basic capability assessments. More comprehensive experiments will be included in an upcoming technical report, which will be released shortly.
@@ 4.1 Datasets
The training data for the Mini-Omni2 model is primarily sourced from five components, as shown in Table 1. (1) Textual Question-Answering Data: Throughout all training stages, whenever the language model weights were unfrozen for training, textual question-answering data was included to maintain the model’s reasoning ability. We used the first 1.5 million question-answer pairs from the Open-Orca dataset. (2) Speech Recognition Data: Speech recognition data was used to continuously maintain the model’s semantic understanding of external spoken input. We primarily utilized the LibriTTS, VCTK, and LibriSpeech datasets. (3) Spoken Question-Answering Data: We did not use a standalone spoken dataset; instead, synthetic data was employed for training. The spoken question-answering data was derived from the Moss-002-sft dataset. (4) Image Question-Answering Data: We used 400,000 samples (caption and instruction) from the ALLaVA-4V dataset. (5) Voice Assistant Data: To make the model’s responses more aligned with the style of voice assistants, we continuously used the VoiceAssistant-400K dataset introduced in Mini-Omni.
@@ 4.2 Training Parameters
The Mini-Omni2 model completed all training steps on eight A100 GPUs. During the adapter training stage, learning rates ranged from 2e-5 to 1e-3, while training the language model used learning rates between 2e-6 and 2e-4. The final fine-tuning was conducted with learning rates ranging from 2e-6 to 2e-5. A cosine scheduler was employed, with 1,500 warm-up steps and a global batch size of 192. Each stage was trained for one epoch using the full dataset. The scale of the vision and audio encoders was described earlier, and the language model used was the Qwen2-0.5B instruct version, without a TTS adapter. All model adapters used Llama-MLP with an intermediate size of 4,864.
@@ 4.3 Data Construction
Spoken Dialogue Data: We used our speech recognition dataset as a random voice timbre library. To ensure robust training, a random sample from this dataset was selected as a voice prompt for the input of all spoken dialogue data, and CosyVoice(Du et al., 2024) was employed for zero-shot speech synthesis. For the output of all question-answering data, the same voice timbre was used.
Interruption Data: First, the noise data is stream-encoded and decoded to simulate real-time streaming input to the model. Then, a random segment of the noise data is extracted. At the end of this segment, a "Stop Omni" phrase is inserted, generated with a random voice timbre in the same manner as the dialogue data. Finally, an additional "tail" of 0-10 seconds is appended to the end of this segment. In terms of labeling, all data before the tail is labeled as "n-irq," while the tail segment is labeled as "irq," indicating that the model should be interrupted by a human.
Identity Data: We replaced all instances of "Moss" in the Moss-002-sft dataset with "Omni" and added 1,000 identity-related response training samples.
(Visual Assistant Data - Coming soon): We found that the currently available visual question answering datasets tend to have very lengthy answers and a very formal tone, which are not suitable for the tone of an assistant. We are building and open-sourcing a new visual QA dataset with a more relaxed and pleasant tone.
@@ 4.4 Experimental Results
In the experimental section, we are comprehensively evaluating the basic capabilities of Mini-Omni2 in speech understanding, image understanding, and text question answering. At the same time, we will test the model’s latency and synthesized speech quality, etc. This is a series of experiments and still requires some time. We will release it in a very quick second version. Currently, we provide the accuracy of Mini-Omni2 in speech recognition to evaluate the model’s speech understanding ability, and we present some practical cases. For model experience and more cases, please follow our github repositories.
@@
Table 2: Comparison of the model’s ASR with the base model used.
@@ 4.5 Case Study
Here we present some real response examples from Mini-Omni2.
Figure 6: Real streaming output examples of Mini-Omni2
@@ 5 Limitations
For the current Mini-Omni2 model, we believe the following aspects are worth exploring and improving: 1. The continued expansion of model size and data scale. Mini-Omni2 aims to train small models with limited resources, and we believe that scaling laws can greatly enhance its capabilities. 2. More powerful encoders and base models. 3. Using multi-modal token-in-token-out, abandoning the pre-trained encoder to enable extensive multimodal pre-training. 4. Control over TTS for voice output (emotion, naturalness, and timbre). 5. A richer mechanism for semantic interruptions.
@@ 6 Conclusion
In this paper, we present Mini-Omni2, a unified multi-modal model with capabilities in text, speech, vision, end-to-end streaming audio output, and duplex interaction. Our goal is to reproduce an open-source GPT-4o model, and to our knowledge, our work is also one of the closest in terms of functionality. We use multiple pretrained encoders as the vision and speech encoders and align them with the language model to extend the modalities.Furthermore, we propose a three-phase modality alignment and expansion training process to achieve the desired capabilities of the model while avoiding an exponential increase in data volume. We also explore a robust method for semantic interruption modeling and introduce our data construction and interruption logic. Finally, we will continue to improve the model’s performance and create multi-modal data resembling vision-speech assistants. All models and datasets will be open-sourced, and we hope Mini-Omni2 can serve as a reference for future research.

### v3

@@ 3.3 Training Strategies
  • Post training In Stage 3, the task of the model is to extend the output modality to include audio response generation. As shown in Figure 5, the model will be trained on all tasks from Stage 1 and Stage 2, with audio token outputs for all question-answering tasks. Additionally, the model will learn interruption mechanism, an algorithm introduced in the next section.
3.4 Duplex Interaction
A real-time conversation model needs to have duplex capability in order to enable more flexible interactions. However, this interruption mechanism should not be a simple VAD (Voice Activity Detection)-based one, but rather a system that can determine whether the user intends to interrupt the model. Additionally, the model’s ability should be highly robust, capable of handling various external situations (e.g., noise, other conversations, and unrelated sounds). We explore this functionality with command-based task, where the model will stop talking immediately when user speaks "Stop Omni". Furthermore, this approach can be naturally extended to incorporate more sophisticated semantic interruption mechanisms through the development of more contextually appropriate interruption datasets.
Background Noise Selection: (1) We randomly utilized a variety of speech recognition samples from the Libri-tts dataset as the original human noise data samples. (2) We employed samples from the MUSAN(Snyder et al., 2015) dataset, which includes music, human voices, white noise, and urban noise.
@@ 3.4 Duplex Interaction
Combining the aforementioned data, the model will receive long sequences of data containing "Stop Omni" phrases amidst various noises. The model will generate two types of state tokens in real time: irq and n-irq, representing the intention of the user to interrupt and not to interrupt, respectively. During inference, when the model output irq token, it will stop the generating process and start to listen the new question. For this task, we use tokens as input to enhance the model’s real-time processing capabilities.
@@ 4 Data and Evaluation
In this section, we introduce the data used for training Mini-Omni2 and present some initial evaluation results. We will provide a more detailed explanation of the composition and construction process of the data for each modality. In the experimental results section, we only showcase a few application cases and basic capability assessments. More comprehensive experiments related text and vision tasks will be updated shortly.
@@ 4.1 Datasets
The training data for the Mini-Omni2 model is primarily sourced from five components, as shown in Table 1. (1) Textual Question-Answering Data: Throughout all training stages, whenever the language model weights were unfrozen for training, textual question-answering data was included to maintain the model’s reasoning ability. We used the first 1.5 million question-answer pairs from the Open-Orca dataset. (2) Speech Recognition Data: Speech recognition data was used to continuously maintain the model’s semantic understanding of external spoken input. We primarily utilized the LibriTTS, VCTK, and Multilingual LibriSpeech datasets. (3) Spoken Question-Answering Data: We did not use a standalone spoken dataset; instead, synthetic data was employed for training. The spoken question-answering data was derived from the Moss-002-sft dataset. (4) Image Question-Answering Data: We used 400,000 samples (caption and instruction) from the ALLaVA-4V dataset. (5) Voice Assistant Data: To make the model’s responses more aligned with the style of voice assistants, we continuously used the VoiceAssistant-400K dataset introduced in Mini-Omni.
@@ 4.2 Training Parameters
The Mini-Omni2 model completed all training steps on eight A100 GPUs. During the adapter training stage, learning rates ranged from 2e-5 to 1e-3, while training the language model used learning rates between 2e-6 and 2e-4. The final fine-tuning was conducted with learning rates ranging from 2e-6 to 2e-5. A cosine scheduler was employed, with 1,500 warm-up steps and a global batch size of 192. Each stage was trained for one epoch using the full dataset. The scale of the vision and audio encoders was described earlier, and the language model used was the Qwen2-0.5B base model. All model adapters used Llama-MLP with an intermediate size of 4,864.
@@ 4.3 Data Construction
Spoken Dialogue Data: We used our speech recognition dataset as a random voice timbre library. To ensure robust training, a random sample from this dataset was selected as a voice prompt for the input of all spoken dialogue data, and CosyVoice(Du et al., 2024) was employed for zero-shot speech synthesis. For the output of all question-answering data, the same voice timbre was used from an internal TTS system.
Interruption Data: First, the noise data is stream-encoded and decoded to simulate real-time streaming input to the model. Then, a random segment of the noise data is extracted. At the end of this segment, a "Stop Omni" phrase is inserted, generated with a random voice timbre in the same manner as the dialogue data. Finally, an additional "tail" of 0-10 seconds is appended to the end of this segment. In terms of labeling, all data before the tail is labeled as "n-irq", while the tail segment is labeled as "irq", indicating that the model should be interrupted.
@@ 4.4 Experimental Results
Currently, we provide the accuracy of Mini-Omni2 in speech recognition to evaluate the model’s speech understanding ability, and we present some practical cases. For model experience and more cases, please follow our github repositories.
@@
Table 2: Comparison of the model’s ASR with the base model used. (* our reproduced evaluation result.)
@@ 4.5 Case Study
Here we present some use cases from Mini-Omni2.
Figure 6: Use cases of Mini-Omni2
@@ 5 Limitations
We believe the following aspects are worth exploring and improving: 1. Scaling of model and data size. Mini-Omni2 aims to train small models with limited resources, and we believe that more data and compute can greatly enhance its capabilities. 2. Improve style control and diversity of audio output (emotion, naturalness, timbre, accent, and singing). 3. A richer mechanism for semantic interruptions.
@@ 6 Conclusion
In this paper, we present Mini-Omni2, a unified multi-modal language model with capabilities in text, speech, vision, end-to-end streaming audio output, and duplex interaction. Our goal is to reproduce an open-source GPT-4o model, and to our best knowledge, our work is also one of the closest in terms of functionality. We use multiple pretrained encoders as the vision and speech encoders and align them with the language model to extend the modalities. Furthermore, we propose a three-phase modality alignment and expansion training process to achieve the desired capabilities of the model. We also explore a robust method for duplex interaction modeling and introduce our data construction and interruption mechanism. All models and datasets will be open-sourced, and we hope Mini-Omni2 can serve as a reference for future research.